OpenStudy (anonymous):

If f (x) = x^2 + 1 and g (x) =1/x - 1. Find the value of (f of g) (10)

OpenStudy (cwrw238):

is g(x) = 1 / (x - 1) ??

OpenStudy (cwrw238):

or (1/x) - 1 ?

OpenStudy (anonymous):

(1/x)-1

OpenStudy (cwrw238):

ok (f of g = ( 1/x - 1 )^2 + 1 = 1/x^2 - 2/x + 2 f of g (10) = 1/100 - 2/10 + 2 = 1/100 - 20/100 + 200/100 = 181/100 = 1.81

OpenStudy (anonymous):

im having a hard time following this..can you use the equation to write it better pls

OpenStudy (cwrw238):

ill try

OpenStudy (cwrw238):

\[f of g = (1divx - 1)^2 + 1\]

OpenStudy (cwrw238):

thats not what i intended

OpenStudy (cwrw238):

\[f of g = (1/x - 1)^2 + 1\]

OpenStudy (cwrw238):

\[= 1/x ^{2} - 2/x + 2\]

OpenStudy (cwrw238):

i dont know how to do more than one line!

OpenStudy (anonymous):

i understand now but how do you get the +2

OpenStudy (cwrw238):

now substitute x = 10 in this exression right expanding the brackets gives 1/x2−2/x+1 then add 1

OpenStudy (cwrw238):

i did two steps in one

OpenStudy (anonymous):

ok thanks

OpenStudy (cwrw238):

1/x2−2/x+1 + 1

OpenStudy (cwrw238):

yw - i'm not very good with the equation thing!

OpenStudy (anonymous):

i dont really mind whether someone uses that or not...it was just extremely hard to follow

OpenStudy (anonymous):

what is f^-1(x) if f(x) = (5x-3)/(2-x) how do you find that

OpenStudy (cwrw238):

ok one way is to let y = (5x-3)/ (2-x) then find x in terms of y then replace y by x in this result this gives f^-1(x) , the inverse of f(x)

OpenStudy (cwrw238):

you need help with that?

OpenStudy (anonymous):

yea

OpenStudy (cwrw238):

ok first cross-multiply y(2-x) = 5x-3 2y - xy = 5x-3 5x + xy = 2y+3 x = (2y+3) / ( y+5) f^-1(x) = (2x+3) / (x+ 5) replacing y's by x's

OpenStudy (cwrw238):

after fourth line i could have typed x(y + 5) = 2y+3 then divide both sides by (y+5)

OpenStudy (anonymous):

thank you very much. i have to leave now i hope i can find you later on

OpenStudy (cwrw238):

yw - sorry - ill be turning in very soon - i'm in uk